# HypeRL: Parameter-Informed Reinforcement Learning for Parametric PDEs

### v1

@@ 4 Numerical Experiments
We validate our proposed approach on two control baselines, namely (i) a parametric Kuramoto-Sivashinsky PDE with distributed in-domain actuators, and (ii) a 2D Navier-Stokes equation with boundary control.
4.1 1D Kuramoto-Sivashinsky equation
@@
The Kuramoto-Sivashinsky (KS) equation is a nonlinear 1D PDE describing pattern and instability in fluid dynamics, plasma physics, and combustion, e.g., the diffusive-thermal instabilities in a laminar flame front [60]. Similarly to [25, 27], we write the KS PDE with state $y(x,t)=y$ and forcing term $u(x,t)=u$ with the addition of a parametric cosine term, breaking the spatial symmetries of the equation and making the search for the optimal control policy more challenging:
$\begin{split}\frac{\partial y}{\partial t}+y\frac{\partial y}{\partial x}+\frac{\partial^{2}y}{\partial x^{2}}+\frac{\partial^{4}y}{\partial x^{4}}+\mu\cos{(\frac{4\pi x}{L})}&=u\,,\\ u&=\sum_{i=1}^{N_{a}}a_{i}\psi(x,m_{i})\,,\\ \psi(x,m_{i})&=\frac{1}{2}\exp(-{(\frac{x-m_{i}}{\sigma})}^{2})\,,\\ \end{split}$ (30)
where $u$ is the control input function with $a_{i}\in[-1,1]$, $\psi(x,c_{i})$ is a Gaussian kernel of mean $c_{i}$ and standard deviation $\sigma=0.8$, $\mu\in[-0.25,0.25]$ is the parameter of interest of the system, and $\mathcal{D}=[0,22]$ is the spatial domain with periodic boundary conditions. Small values of $\mu$ generates chaotic solutions, while large values of $\mu$ periodic solutions, making it a very challenging test case and requiring strong generalization capabilities of the control policies. To numerically solve the PDE, we discretize the spatial domain with $N_{x}=64$. We assume to have $N_{a}=8$ equally-spaced actuators. The state of the agent is set equal to the (discretized) PDE state (although we are not limited to that). Similarly to [25, 27], we utilize the reward function (see Equation (25)) with $\bm{y}_{\text{ref}}=\bm{0}$, $\bm{u}_{ref}=\bm{0}$, and $\alpha=0.1$. The choice of $\alpha=0.1$ is dictated by the need for balancing the contribution of the state cost $c_{1}$ and the action cost $c_{2}$. In particular, in our experiments we prioritize steering the system to the reference state over the minimization of the injected energy. We train the control policies by randomly sampling a value of the parameter $\mu$ at the beginning of each training episode $\in[-0.225,-0.2,-0.175,-0.15,-0.125,-0.1,-0.075,-0.05,-0.025,0.0,0.025,0.05,0.075,0.1,0.125,0.15,\allowbreak 0.175,0.2,0.225]$. To test the generalization abilities of the policies, we evaluate the agents on unseen and randomly-sampled parameters $\in[-0.25,0.25]$. We compare our two variants of HypeRL TD3 agent (see Section 3) with (i) the standard TD3 agent with state $\bm{y}_{k}$ augmented by the parameter $\mu$, and (ii) the TD3 algorithms without access to the parameter $\mu$.
In Table 1 and 2, we show the training and evaluation rewards collected by the different agents over 5 different random seeds, where we highlight in bold the highest scores in three different phases of the training/evaluation. We consider the average cumulative reward of 5 independent runs over the whole training/evaluation episodes, the average cumulative rewards collected after 500/10 episodes until the end, and the average cumulative rewards collected after 1000/20 episodes until the end. The evaluation is performed by randomly-sampling 10 different instances of the parameter $\mu$ and then record the agents performance.
Cumulative reward mean $\pm$ std mean $\pm$ std ($>500$ ep.) mean $\pm$ std ($>1000$ ep.)
HypeRL-TD3 $-160.06\pm 81.35$ $-130.56\pm 53.71$ $\bm{-112.66}\pm 48.29$
HypeRL-TD3 ($\mu$ only) $\bm{-149.09}\pm 79.01$ $\bm{-129.65}\pm 69.19$ $-118.21\pm 76.41$
TD3 $-186.59\pm 72.56$ $-168.28\pm 61.39$ $-148.12\pm 42.04$
TD3 (no $\mu$) $-197.98\pm 97.52$ $-165.62\pm 63.06$ $-150.08\pm 49.58$
Table 1: Mean and standard deviation of the cumulative reward over training collected by the different algorithms. The results report the average performance over 5 different random seeds.
Cumulative reward mean $\pm$ std mean $\pm$ std ($>10$ ep.) mean $\pm$ std ($>20$ ep.)
HypeRL-TD3 $-163.68\pm 62.35$ $\bm{-136.95}\pm 37.63$ $\bm{-114.28}\pm 32.70$
HypeRL-TD3 ($\mu$ only) $\bm{-150.09}\pm 37.40$ $-138.21\pm 34.84$ $-122.98\pm 37.88$
TD3 $-189.67\pm 43.88$ $-174.59\pm 30.75$ $-154.58\pm 17.97$
TD3 (no $\mu$) $-193.29\pm 77.11$ $-165.16\pm 32.39$ $-164.49\pm 42.13$
Table 2: Mean and standard deviation of the cumulative reward over evaluation collected by the different algorithms. The results report the average performance over 5 different random seeds.
The results show that the hypernetwork-based algorithms, namely HypeRL-TD3 and HypeRL-TD3 ($\mu$ only), outperform the TD3 agents not only in later stages of the training but also in early ones, showing that the way the information about the PDE parameter is encoded is crucial for sample efficiency and generalization of the RL algorithms. HypeRL-TD3 ($\mu$ only) is the algorithm capable of achieving the highest average training rewards, followed by HypeRL-TD3, TD3, and TD3 (no $\mu$). However, if we only consider later stages of the training HypeRL-TD3 is the best performing agent. We see similar trends in the evaluation results.
@@
In Figure 3 and Figure 10, we show the mean and the standard deviation of the cumulative reward, state cost, and action cost during training and evaluation, respectively. It is possible to notice the superior performance of the hypernetwork-based agents in early and later stages of the training and during evaluation with unseen values of the parameter $\mu$.
@@
Figure 3: Training and evaluation results. The solid line represents the mean and the shaded area the standard deviation over 5 different random seeds.
Additionally, due to the stochastic nature of the experiments and to improve the reliability of the results, we compute and show in Figure 4 the $95\%$ confidence intervals, as suggested in [61].
Figure 4: $95\%$ confidence intervals of the training and evaluation reward for the different phases of training and evaluation. This metric, suggested in [61], allows to assess the reliability of the results accounting for the stochastic nature of the RL experiments.
The $95\%$ confidence intervals are aligned and consistent with the average training and evaluation results and again highlights the higher sample efficiency and generalization capabilities of the hypeRL agents.
Eventually, in Figure 5 we show examples of controlled solution of the KS PDE. The KS PDE is evolved for 100 time-steps before turning the controllers ON. In particular, we highlights two different test cases: (a) interpolation regime with $\mu=0.062$ (unseen but withing the training range), and (b) extrapolation regime with $\mu=0.25$ (unseen and outside the training range). We also report the state and action costs of the solutions after enabling the controllers.
Figure 5: Controlled solutions for $\mu=-0.062$ and $\mu=0.25$.
The hypeRL agents are capable of quickly steering the state of the KS PDE very close to the reference value with very little control effort in interpolation and extrapolation regimes, even when the PDE dynamics is very different. On the other side, the TD3 agents (with and without access to the PDE parameter $\mu$) struggle to achieve good state tracking, especially in extrapolation regimes. In both interpolation and extrapolation the hypeRL agents improve the results with respect to the TD3 agents, showing that learning the main networks weights’ with a PDE-parameter dependent hypernetwork is the key ingredient for improving generalization and sample efficiency of RL algorithms.
4.2 2D Navier-Stokes equations
As a second test case, we consider a 2D parametric incompressible Navier-Stokes (NS) equations. The NS equations are in fluid dynamics with applications in different fields of engineering and science [62]:
$\begin{split}\nabla\cdot\bm{y}&=0\\ \frac{\partial\bm{y}}{\partial t}+\bm{y}\cdot\nabla\bm{y}&=-\frac{1}{\rho}\nabla p+\mu\nabla^{2}\bm{y}\,,\end{split}$ (31)
where $\bm{y}=(y_{1},y_{2}):\mathcal{X}\times[t_{0},t_{f}]\to\mathbb{R}^{2}$ represents the 2D velocity field, $\mathcal{X}=[0,1]\times[0,1]$ is the spatial domain, $\mu$ is the kinematic viscosity, $\rho$ is the fluid density, and $p$ is the pressure field. To numerically solve the PDE, we discretize the spatial domain in $21\times 21\times 2=882$ dimensions. The control is applied from one of the boundaries of the domain $\bm{u}(x_{1}=x,x_{2}=0,t)=u_{k},\forall x\in[0,1]$, where $x_{1},x_{2}$ represent the spatial variables in the 2D domain. We use Dirichlet conditions for all the other boundaries. Similarly to the KS, we utilize the reward function in Equation (25) where $\bm{y}_{\text{ref}}$ is obtained when applying the controls $u=3-5t$ for $t\in[t_{0}=0,t_{f}=T]$, $u_{\text{ref}}=2.0$, $T=0.2$, and $\alpha=0.01$. It is worth mentioning that the reference solution is the one provided by PDEControlGym with a default and fixed value of the kinematic viscosity $\mu$ and it is kept the same for all the difference instances of the parameter. The initial velocity field and the initial pressure are sampled from a uniform distribution $\in(-5,5)$.
We extend the implementation of PDEControlGym [63] to include a variations of the kinematic viscosity $\mu$, our parameter of interest. Changing the kinematic viscosity is an effective way to change the Reynolds number ($Re\propto 1/\mu$). In particular, we vary $\mu\in[0.01,0.025,0.05,0.075,0.1]$ during training, while we evaluate the different agents for $\mu\in[0.009,0.12]$. The agents observe the (discretized) velocity field $\bm{y}_{k}$ at different time-steps and predict the 1D control action $u_{k}$. We use the same main network and hypernetwork architectures of the KS case. However, due to the high-dimensionality of the state ($21\times 21\times 2$), we utilize a convolutional encoder to reduce the state dimensionality to $20$ dimensions before feeding it to the RL networks. We discuss the architecture of the convolutional encoder in Appendix A. Again, we compare our two variants of HypeRL TD3 agent (see Section 3) with (i) the standard TD3 agent with state $\bm{y}_{k}$ augmented by the parameter $\mu$, and (ii) the TD3 algorithms without access to the parameter $\mu$.
@@
In Table 3 and 4, we show the training and evaluation rewards collected by the different agents, where we highlight in bold the highest scores in three different phases of the training/evaluation. We consider the average cumulative reward over the whole training/evaluation episodes, the cumulative rewards collected after 250/40 episodes until the end, and the cumulative rewards collected after 450/80 episodes until the end.
Cumulative reward mean $\pm$ std mean $\pm$ std ($>250$ ep.) mean $\pm$ std ($>450$ ep.)
HypeRL-TD3 $\bm{-13.50}\pm 4.00$ $\bm{-13.47}\pm 4.10$ $\bm{-13.07}\pm 3.85$
HypeRL-TD3 ($\mu$ only) $-13.64\pm 4.02$ $-13.57\pm 4.09$ $-13.16\pm 3.83$
TD3 $-13.80\pm 4.10$ $-13.68\pm 4.18$ $-13.24\pm 3.94$
TD3 (no $\mu$) $-13.97\pm 4.04$ $-13.88\pm 4.13$ $-13.41\pm 3.84$
@@
Table 3: Mean and standard deviation of the cumulative reward over training collected by the different algorithms.
Cumulative reward mean $\pm$ std mean $\pm$ std ($>40$ ep.) mean $\pm$ std ($>80$ ep.)
HypeRL-TD3 $\bm{-10.21}\pm 1.62$ $-9.69\pm 0.89$ $-9.45\pm 0.88$
HypeRL-TD3 ($\mu$ only) $-10.37\pm 1.88$ $-9.70\pm 0.93$ $-9.41\pm 0.02$
TD3 $-10.58\pm 2.36$ $\bm{-9.68}\pm 0.90$ $\bm{-9.40}\pm 0.87$
TD3 (no $\mu$) $-10.77\pm 2.24$ $-9.90\pm 0.91$ $-9.62\pm 0.89$
Table 4: Mean and standard deviation of the cumulative reward over evaluation collected by the different algorithms.
Even in the NS case, the HypeRL agents still outperform the TD3 agents. However, because the control is only applied on one boundary, the system is less "controllable" than the KS PDE and the difference in performance among the different agents is smaller. For example, if we only consider the later stages of the training, the TD3 agent is capable of achieving similar performance to the HypeRL agents in evaluation, while still inferior in early training/evaluation stages.
In Figure 6, 11, and 7, we show average rewards, state and action costs, and $95\%$ confidence bounds of the rewards collected by the different agents.

### v2

@@ 3.2 HypeRL-TD3
3.2.1 Neural Network Architectures
The main policy network $\pi(\bm{z}_{k};\bm{\theta}_{\pi})$ is composed of one input layer of dimension equal to the agent’s state dimension $\bm{z}_{k}\in\mathbb{R}^{|\bm{y}|+|\bm{\mu}|}$, one hidden layer with $256$ neurons, and an output layer of dimension equal to the control action dimension $|\bm{u}_{k}|\in\mathbb{R}^{|\bm{u}|}$. In particular, we can write the main policy network as:
[…]
where $\bm{g}^{(1)},W^{(1)},\bm{b}^{(1)}$ and $\bm{g}^{(2)},W^{(2)},\bm{b}^{(2)}$ are respectively the gain vectors, the weights matrices, and the biases of the two layers of the main networks that are learned via the hypernetwork $h(\bm{z}_{k};\bm{\theta}_{h_{\pi}})$, i.e., the hypernetworks outputs, $\bm{\theta}_{\pi}=[\bm{g}^{(1)},W^{(1)},\bm{b}^{(1)},\bm{g}^{(2)},W^{(2)},\bm{b}^{(2)}]$, and $\odot$ indicates the element-wise product. This type of architecture is motivated by the work in [40] and it is depicted in Figure 2.
Figure 2: Hyper-policy architecture.
Similarly to the policy, the main value function networks $Q_{1}(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q_{1}})$ and $Q_{2}(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q_{2}})$ are composed of one input layer of dimension $|\bm{z}_{k}|+|\bm{u}_{k}|$, one hidden layer with $256$ neurons, and an output layer of dimension 1. By indicating with $\bm{\upsilon}_{k}=[\bm{z_{k},\bm{u}_{k}}]$, we can write:
[…]
The hypernetworks use the same architecture and weight initialization proposed in [49]. Each hypernetwork is composed of three blocks that map $\bm{z}_{k}$ to $\bm{g}^{(1)},W^{(1)},\bm{b}^{(1)}$ and $\bm{g}^{(2)},W^{(2)},\bm{b}^{(2)}$. Each block is composed of a linear layer and two residual blocks and each residual block is composed of two linear layers with ReLU activation and a skip connection.
@@ 4 Numerical Experiments
We validate our proposed approach on two control baselines, namely (i) the stabilization of a parametric Kuramoto-Sivashinsky equation with distributed in-domain actuators and sensors, and (ii) a 2D navigation problem in a gyre flow, where the goal is to control the velocity of particle to reach arbitrary targets.
We compare the HypeRL-TD3 agent with:
  1. 1.
    TD3. The algorithm uses the default and widely-used architecture for policy and value functions [19] composed of an input layer, two hidden layers with $256$ neurons each and an output layer. Analogously to HypeRL-TD3, the input of the policy networks is $\bm{z}_{k}$ and the input of the value function networks is $\bm{z}_{k},\bm{u}_{k}]$.
  2. 2.
    TD3 (Huber). This is a variant of TD3 where the Huber loss replaces the mean-squared error loss to train the critic networks to match the objective used by HypeRL-TD3.
  3. 3.
    TD3 (2 layers). This a variant of TD3 where a single hidden layer is used to match the HypeRL-TD3 main policy and value function networks.
  4. 4.
    TD3 (no $\bm{\mu}$). We remove the knowledge of the parameter vector $\bm{\mu}$ to TD3 to show the need for making the agents parameter-informed in parametric control problems.
In all our numerical experiments, each agent is trained for $2000$ episodes. The evaluation occurs every $200$ training episodes and we apply a warmup of $100$ episodes in which the agents’ actions are sampled from a uniform distribution $\in\mathcal{U}(-\bm{u}_{\min},\bm{u}_{\max})$. The evaluation metric is the average cumulative reward of $10$ evaluation episodes in which the optimal deterministic policy is deployed.
4.1 Stabilization of a 1D Kuramoto-Sivashinsky Equation
@@
The Kuramoto-Sivashinsky (KS) equation is a nonlinear 1D PDE describing pattern and instability in fluid dynamics, plasma physics, and combustion, e.g., the diffusive-thermal instabilities in a laminar flame front [31]. Similarly to [44, 8], we write the KS PDE with state $s(x,t)$ and forcing term $u(x,t)$ with the addition of a parametric cosine term, breaking the spatial symmetries of the equation and making the search for the optimal control policy more challenging:
[…]
where $u$ is the time-dependent control input function with $a_{i}(t)\in[-1,1]$, $\psi(x,c_{i})$ is a Gaussian kernel of mean $c_{i}$ and standard deviation $\sigma=0.8$, $\nu\in[-0.25,0.25]$ is the parameter of interest of the system, and $\mathcal{D}=[0,22]$ is the spatial domain with periodic boundary conditions. We numerically solve the KS equation for $T=300$s with a $dt=0.1$ and we discretize the spatial domain with $N_{x}=64$, leading to $s(x,t_{k})\mapsto\bm{y}_{k}\in\mathbb{R}^{N_{x}}$ for each time instance $t_{k}$. In Figure 3 we show examples of solution for different values of the parameter $\nu$. It is worth highlighting that small variations of $\nu$ lead to vastly different solutions.
Figure 3: Solutions of the uncontrolled KS equation for different values of $\nu$, namely, $\nu=0$ (top left), $\nu=0.2$ (top right), $\nu=-0.25$ (bottom left), and $\nu=-0.1$ (bottom right).
We assume to have $N_{a}=8$ equally spaced actuators. We evolve the KS equation for $100$ timesteps before activating the controller. Similar settings were used in [44, 8]. However, instead of looking for a controller that solely stabilizes the KS state to the zero reference, i.e., $\bm{y}_{\text{ref}}=\bm{0}$, we consider two more challenging control problems:
  1. 1.
    stabilization of the state of a KS equation to an arbitrary reference $\in\mathcal{U}(-3.0,3.0)$ with $\nu=0.0$, and
  2. 2.
    stabilization of the state of a parametric KS equation to an arbitrary reference $\in\mathcal{U}(-3.0,3.0)$ with $\nu\in\mathcal{U}(-0.25,0.25)$  ,
where $\mathcal{U}$ indicates the uniform distribution. In both cases, the agent’s state $\bm{z}_{k}=[\bm{y}_{k},\bm{\mu}]$, where we indicate with $\bm{\mu}=[\bm{y}_{\text{ref}},\nu]$.
The reward function corresponds to the fully-discretized (in space and time) counterpart of the running cost in Equation (3) with opposite sign. Our goal is to steer the measured state $\bm{y}_{k}$ to a reference value $\bm{y}_{\text{ref}}$ with minimal control effort. Therefore, we utilize the following reward function:
[…]
where $\bm{y}_{\text{ref}}$ indicates the reference values for the state, and $\alpha=0.1$ is a scalar positive coefficient balancing the contribution of the two terms.
4.1.1 Stabilization of the State of a KS Equation to an Arbitrary Reference
In this first example, the goal is to steer the state of a KS equation to an arbitrary reference that is randomly sampled from a uniform distribution in $\in\mathcal{U}(-3.0,3.0)$ at the beginning of each training and evaluation episode.
@@ 4.1.1 Stabilization of the State of a KS Equation to an Arbitrary Reference
In Figure 4, we show the mean and the standard deviation of the cumulative reward during training and evaluation, respectively. It is possible to notice the superior performance of the hypernetwork-based agent in early and later stages of the training and during evaluation.
@@
Figure 4: Stabilization of the state of a KS equation to an arbitrary reference – training and evaluation results. The solid line represents the mean and the shaded area the minimum and maximum values observed over 5 different random seeds.
@@
In Table 1, we show the training and evaluation rewards collected by the different agents over 5 different random seeds, where we highlight in bold the highest scores of the training and evaluation. In the calculation of these scores, we remove for every agent the rewards collected during $100$ warmup episodes as common to each agent.
@@
Table 1: Mean and standard deviation of the cumulative reward over training (left) and evaluation (right) collected by the different algorithms. The results report the average performance over 5 different random seeds.
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-180.41\pm 62.97}$
TD3 $313.17\pm 116.74$
TD3 (Huber) $263.78\pm 93.49$
TD3 (2 layers) $344.13\pm 119.91$
TD3 (no $\bm{\mu}$) $12522.60\pm 25874.76$
Average cumulative reward mean $\pm$ std
HypeRL-TD3 $\bm{-175.33\pm 20.85}$
TD3 $304.27\pm 63.11$
TD3 (Huber) $249.76\pm 37.31$
TD3 (2 layers) $344.52\pm 68.47$
TD3 (no $\bm{\mu}$) $17473.19\pm 31416.09$
The results show that the HypeRL-TD3 outperforms the TD3 agents, showing that the way state and parametric information is encoded are crucial ingredients for sample efficiency and generalization of the RL algorithms.
In Figure 5, we show an example of controlled solution for $\bm{y}_{\text{ref}}=-\bm{1}$, where the first row depicts the evolution of the state, the second one of the control, and the third one of the state-tracking error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$. Additional results, including the performance of the TD3 agent without access to $\bm{\mu}$, and state and action costs over training and evaluation, as defined in Equation (15), are in Appendix B.1.
Figure 5: Controlled solutions of the KS equation when using the different agents. The first row shows the state, the second the control, and the third one the error $|\bm{y}_{k}-\bm{y}_{\text{ref}}|$.
4.1.2 Stabilization of the State of a Parametric KS Equation to an Arbitrary Reference
In this second example, the goal is to steer the state of a parametric KS equation to an arbitrary reference that is randomly sampled from a uniform distribution in $\in\mathcal{U}(-3.0,3.0)$ at the beginning of each training and evaluation episode together with the parameter $\nu\in\mathcal{U}(-0.25,0.25)$.
In Figure 6, we show the mean and the standard deviation of the cumulative reward during training and evaluation, respectively. It is possible to notice the superior performance of the hypernetwork-based agent in early and later stages of the training and during evaluation.